# Working Backwards: Learning to Place by Picking

### 04 Dec 2023 v1

@@ Abstract
We present Learning to Place by Picking (LPP), a method capable of autonomously collecting demonstrations for a family of placing tasks in which objects must be manipulated to specific locations. With LPP, we approach the learning of robotic object placement policies by reversing the grasping process and exploiting the inherent symmetry of the pick and place problems. Specifically, we obtain placing demonstrations from a set of grasp sequences of objects that are initially located at their target placement locations. Our system is capable of collecting hundreds of demonstrations without human intervention by using a combination of tactile sensing and compliant control for grasps. We train a policy directly from visual observations through behaviour cloning, using the autonomously-collected demonstrations. By doing so, the policy can generalize to object placement scenarios outside of the training environment without privileged information (e.g., placing a plate picked up from a table and not at the original placement location). We validate our approach on home robotic scenarios that include dishwasher loading and table setting. Our approach yields robotic placing policies that outperform policies trained with kinesthetic teaching, both in terms of performance and data efficiency, while requiring no human supervision.
@@ I Introduction
Object placing is an essential everyday task that autonomous robot manipulators will need to perform reliably. Many household chores, such as setting the table and loading a dishwasher, for example, require careful placing. Robotic placement involves contact-rich interactions and typically necessitates sensory feedback to determine the state of the in-hand object and the environment. Designing object placement policies is non-trivial and can involve solving a number of complicated subtasks, including object tracking, scene understanding, motion planning, and control.
Imitation learning (IL) provides an appealing, simple, and practical alternative by learning end-to-end control policies from expert demonstrations [1, 2]. The real world is complex and demonstration data should ideally include diverse examples to account for varying factors such as background, lighting conditions, and object types. Some recently-developed robotic systems [3, 4] have demonstrated the use of IL with a large amount of demonstration data to effectively solve a variety of tasks. A major bottleneck for IL-based approaches in these data regimes is the human time and effort required to collect expert demonstrations.
We propose a novel approach named Learning to Place by Picking (LPP) that facilitates and automates the collection of expert demonstrations for a large subset of placing problems by taking advantage of powerful grasp planners [5]. LABEL:fig:title provides a visual summary of our approach. LPP is a simple self-supervised pipeline to autonomously collect expert placing demonstrations by leveraging the cyclical nature and inherent symmetry of the pick and place problems. Given an environment with objects initially positioned at their target place locations, we alternate between picking (i.e., grasping and retrieving) and placing, by time-reversing the retrieval trajectory. In this way, we are able to generate place demonstration data in a self-supervised manner—the picking phase provides the training supervision for the place task. Finally, we learn a placing policy based only on visual observations using the demonstrations collected in the training environment. Crucially, this allows our reactive policy to be used for object placement outside of the training environment and in scenarios where the objects are not initially placed at their target locations (i.e., without privileged information). Our main contributions are:
@@
    LPP, a self-supervised data collection method for 6-DOF robotic object placements,
@@
    a flexible grasp planning pipeline that enables a user to specify what to grasp in a scene,
@@
    a tactile-driven regrasping and compliant control strategy to achieve uninterrupted pick and place data collection without human intervention, and
@@
    real-world experimental validation of LPP by using the collected demonstrations to train a vision-based policy capable of placing multiple plates in a dishrack and of setting a table.
@@ II Related Work
In this section, we review previous approaches to object placing with manipulators and other work that generally leverage the idea of “working backwards”.
II-A Robotic Manipulator Placing
Pick and place is a fundamental problem in robotic manipulation. A significant amount of research has focused on the picking problem [6, 7, 8, 9], while placing is relatively less explored.
Previous approaches have explored the use of motion planning combined with place-specific objectives [10, 11]. For example, in [10], the authors demonstrate a hierarchical sampling-based motion planner that first finds poses which satisfy constraints (e.g., reachability) and then performs a local optimization for user-specified objectives (e.g., clearance). The place problem has also been studied from the perspective of geometric modelling [12], where the planar surfaces of a discretized object model are matched to planar patches in the environment. Importantly, 3D graphic models of the objects being placed as well as the environment are assumed to be available ahead of time. Other works have looked into combining learning with motion planning to operate directly on point clouds, thereby relaxing the assumption of known object and environment models [13, 14].
Closer to our approach, other works have trained end-to-end place policies using reinforcement learning (RL) [15] and imitation learning (IL) [16, 17]. In [16], the authors demonstrate an inverse reinforcement learning (IRL) technique capable of simultaneously learning a cost function and a control policy from human kinesthetic demonstrations of dish placing. The main contribution and novelty of our work is a data collection technique for robotic placing tasks that can be used in combination with any existing methods that require expert demonstrations.
@@ II-B Working Backwards
The general idea of working backwards or time-reversal has been exploited in multiple contexts including generative modelling [18, 19] and in learning visual predictive models for control [20, 21, 22]. Our self-supervised data collection method, which extends the concept of working backwards to the problem of robotic placement, is most similar in spirit to [23, 24, 25]. These methods consider objects that are assumed to be in their desired place configuration during data collection. The manipulator can then pick an object and move back to the initial goal location to generate data. However, we focus on placement as opposed to insertion as our task. Insertion policies generally only reason about local interactions, while placement involves reasoning over a larger scene with the potential for multiple objects and place poses to be involved in the task. Furthermore, unlike prior work [23, 24, 25], our system does not require a human to guide the manipulator to the object and instead leverages tools from grasp planning. Our approach also resembles [26], where time-reversed trajectories are used to generate assembly data from disassembly data in a self-supervised manner. However, we focus on generating demonstration data for a closed-loop reactive policy that uses a 6-DOF pose-based action representation for placement. In contrast, the self-supervised data collection procedure introduced in [26] uses a 2D open loop pick-and-place action primitive.
III Learning to Place by Picking
We first present an overview of our self-supervised data collection method, LPP, in Section III-A and describe a method for noise augmentation during data collection in Section III-B. We then provide details on the procedure used to train our IL policy in Section III-C.
@@ III-A Self-supervised Data Collection
Figure 1: The four steps involved in LPP, our autonomous demonstration data collection process for placing. We (a) generate grasps with an off-the-shelf grasp planner [5], (b) compliantly grasp the object to apply minimal forces to the environment while ensuring a stable grasp with tactile sensing, (c) retrieve the object with rotational compliance while storing the trajectory, and (d) generate placement demonstration data by rolling out the reversed grasp trajectories while storing the observations and actions.
LPP consists of a cycle with four main phases: (1) grasp planning, (2) grasping, (3) retrieving and (4) placing by reversing; these steps are visualized in Figure 1. During training, the objects of interest are initially in their goal positions.
@@ III-A1 Grasp Planning
Figure 2: The steps involved in our language-driven grasp planning pipeline. We use (a) Grounding-Dino [27] for object bounding box detection based on user provided texts and (b) Segment Anything [28] on the cropped images to produce object specific masks. We then use (c) Contact-graspnet [5] for grasp generation on only the segmented objects. Top: We specify “green plate” and “blue plate” as the targets for data collection. Bottom: We specify “bowl” and “mug” as the targets for data collection.
At the start of the data collection process, the manipulator moves to a pre-determined pose where the camera has a clear view of the objects of interest in the environment, as shown in Figure 1. We use the grasp planner Contact-GraspNet [5] to generate $L$ candidate grasp poses $\mathbf{T}_{g}=\{\mathbf{T}_{g,i}\}_{i=0}^{L}$, where $\mathbf{T}_{g,i}\in SE(3)$ for all $i\in\{0,...,L\}$. The user then provides text-based descriptions of the objects of interest (e.g., “plates”, “green plate”, “cup” or, more generally, “objects”). Given the user-provided texts, we use Grounding-Dino [27] to find the object bounding boxes. Finally, we run Segment Anything [28] on each of the previously generated image crops from the bounding boxes to generate per-object masks, which Contact-GraspNet can then use to filter grasps that are not of objects of interest: $\bar{\mathbf{T}}_{g}=\{\mathbf{T}_{g,i}\}_{i=0}^{K}$, where $K\leq L$ and $\bar{\mathbf{T}}_{g}\subseteq\mathbf{T}_{g}$. We visualize an example of this process in Figure 2 for two different scenes. After pruning, we randomly select a single grasp pose, $\mathbf{T}_{g,k}$, from $\bar{\mathbf{T}}_{g}$. To approach the object of interest without collisions, we include a pregrasp pose, $\mathbf{T}^{\text{pre}}_{g,k}$, that is defined with a translational offset above the grasp pose $\mathbf{T}_{g,k}$.
@@ III-A2 Grasping
Given the computed pregrasp and grasp poses, we move to each pose sequentially by linearly and spherically linearly interpolating to produce a smooth trajectory. An example of a grasp sequence is shown in Figure 1. During grasping, two modules play a critical role in the robustness of LPP: compliant control for grasping (CCG) and tactile regrasping (TR).
@@
Compliant Control for Grasping We use a cartesian impedance controller for manipulation. This choice allows us to set the level of compliance of the manipulator at various stages as needed. Before closing the gripper to grasp, we set the translational and rotational stiffness of the manipulator’s controller to minimal values. The low compliance of the robot arm then naturally allows the gripper to comply with the contact constraints and minimize its applied forces against the environment.
Tactile Regrasping Our method assumes that the grasps that we perform are stable, that is, the object does not move or slip once grasped. To mitigate this issue, we use tactile sensing to preemptively detect if a grasp is stable and perform a regrasp if needed.
III-A3 Retrieving
We begin the retrieval phase when a valid stable grasp is detected. An image of the retrieval procedure is displayed in Figure 1. We store the pose of the end effector (EE) as well as the respective timestamp throughout the retrieval process. The manipulator is set to be compliant along the rotational axes of the EE when commanded to move to the pregrasp pose. Doing so allows the manipulator to better comply to any unexpected interactions with the environment. Once at the pregrasp pose, we randomly sample a clearance pose $\mathbf{T}_{c,k}$ centered around a fixed pose above the scene. That is, given a fixed clearance pose, we add random translations sampled from $\mathcal{N}(0,\,\sigma_{tr}^{2})$. We use $\sigma_{tr}$ = 2.5 $\mathrm{cm}$ during data collection. At the end of the retrieval phase, we have an expert retrieval trajectory of length $M$ from grasp pose to clearance pose via a pregrasp pose:
where $\mathbf{T}_{0}=\mathbf{T}_{g,k}$, $\mathbf{T}_{M}=\mathbf{T}_{c,k}$ and $t_{0},...,t_{M}$ are the respective timestamps of the trajectory starting with $t_{0}$ = 0. To match the desired control frequency of our policy, we sample states $\Delta t$ apart to generate a more sparse trajectory $\bar{\tau}^{\text{expert}}_{r}$ of length $\bar{M}<M$. Our visual policies operate at 5 $\mathrm{h}\mathrm{z}$, so we set $\Delta t=$ 0.20 $\mathrm{s}$.
III-A4 Placing by reversing
We reverse the sparse retrieval trajectory $\bar{\tau}^{\text{expert}}_{r}$ to extract a desired place trajectory:
$\bar{\tau}^{\text{expert}}_{p}=(\mathbf{T}_{\Delta t_{\bar{M}}},...,\mathbf{T}_{\Delta t_{0}}).$ (2)
$\mathbf{T}_{\Delta\text{cmd},i}=\mathbf{T}^{-1}_{\Delta t_{\bar{M}-i}}\mathbf{T}_{\Delta t_{\bar{M}-i-1}},$ (3)
where $i\in\{1,...,\bar{M}\}$ and $\mathbf{T}_{\Delta\text{cmd},i}\in SE(3)$ is a relative change in EE pose. In practice, we use six-dimensional vectors $[\mathbf{t},\boldsymbol{\theta}]^{T}$ to represent poses, where $\mathbf{t}\in\mathbb{R}^{3}$ is a translation vector and $\boldsymbol{\theta}\in\mathbb{R}^{3}$ is a rotation vector. The rotation vector is defined as $\boldsymbol{\theta}=\theta\mathbf{e}$, where $\theta\in\mathbb{R}^{1}$ is the rotation angle and $\mathbf{e}\in\mathbb{R}^{3}$ is the rotation axis. We define the expert action as $\mathbf{a}_{i}=(\mathbf{a}_{\Delta\text{cmd},i},a_{\text{gripper},i})$ where $\mathbf{a}_{\Delta\text{cmd},i}=[\mathbf{t}_{i},\boldsymbol{\theta}_{i}]^{T}$ is the six-dimensional vector representation of the pose $\mathbf{T}_{\Delta\text{cmd},i}$ and $a_{\text{gripper},i}\in\{0,1\}$ is a binary variable describing the gripper command (open or close). To collect an episode of expert data, we perform a rollout of the downsampled place commands, $\{\mathbf{a}_{i}\}_{i=0}^{\bar{M}}$, and store the sequences of observations and actions as an expert trajectory. We use RGB images from the wrist camera to form the robot’s observation space $\mathbf{o}_{i}$, where $\mathbf{o}_{i}\in\mathbb{R}^{w\times h\times 3}$. We downsample the images to a resolution of $128\times 128$. During the place demonstration, the robot keeps its gripper closed (i.e., the robot uses the following gripper command $a_{\text{gripper},i}=1$ for all $i\in\{0,...,\bar{M}\}$). Finally, at the end of the trajectory, we command the robot to keep the EE in its current pose and to open the gripper for N time steps (i.e. $\mathbf{a}_{\Delta\text{cmd},j}=[0,0,0,0,0,0]^{T}$ and $a_{\text{gripper},j}=0$ for all $j\in\{\bar{M}+1,...,\bar{M}+N\}$). We use $N=5$, which gives the robot an adequate amount of time to open the gripper and release the object. A single expert demonstration or trajectory is then defined as a set of training tuples:
$\tau^{expert}=\{\mathbf{o}_{t},\mathbf{a}_{t},\mathbf{o}_{t+1}\}_{t=0}^{T},$ (4)
with a total length of $T=\bar{M}+N$.
III-B Noise-Augmented Data Collection
To increase the coverage of our demonstrations and train more robust policies, we perturb the first $75\%$ of the poses with noise. We draw inspiration from previous works that have investigated the use of noise injection for both the policy observations [29, 30], and actions or controls [31]. We call this variant noise-augmented data collection. We perturb the translation of the pose $\mathbf{t}$ as follows:
$\mathbf{t}_{perturbed}=\mathbf{t}+\delta\mathbf{t},$ (5)
where $\delta\mathbf{t}\in\mathbb{R}^{3}$ is isotropic Gaussian noise and each dimension of $\delta\mathbf{t}$ is sampled from $\mathcal{N}(0,\sigma_{t})$. Similarly, we perturb the rotation as follows:
$\boldsymbol{\theta}_{perturbed}=(\theta+\delta\theta)(\mathbf{e}+\delta\mathbf{e}),$ (6)
where $\delta\mathbf{e}\in\mathbb{R}^{3}$ is isotropic Gaussian noise and each dimension of $\delta\mathbf{e}$ is sampled from $\mathcal{N}(0,\sigma_{e})$, and $\delta\theta$ is sampled from $\mathcal{N}(0,\sigma_{\theta})$. We find $\sigma_{t}=0.5$\mathrm{c}\mathrm{m}$$, $\sigma_{e}=0.5$\mathrm{c}\mathrm{m}$$ and $\sigma_{\theta}=$$$ 0.5  ° to be reasonable values. We evaluate the effects of this augmentation on policy performance in Section IV-B
III-C Policy Learning
We learn a policy $\pi_{\boldsymbol{\phi}}(\mathbf{a}_{t}|\mathbf{o}_{t})$ parameterized by $\boldsymbol{\phi}$ using data generated by LPP. We use behavioral cloning (BC) [32] to train our policy with a likelihood-based loss or objective function. Given a dataset of $N_{train}$ demonstration trajectories, our loss function is then:
The main backbone of our policy network consists of a convolutional neural network (CNN) based on the ResNet18 architecture [33], followed by a multilayer perceptron (MLP) layer that maps the CNN output into the parameters of the distribution of actions. In this work, we consider two representations for the action distribution $\pi_{\phi}(\mathbf{a}|\mathbf{o})$: a unimodal Gaussian and a multimodal mixture of Gaussians. In both cases, during inference or control, we use a low-noise evaluation scheme, as done in [34, 35, 36], by setting the standard deviation of the Gaussian distributions to be the arbitrarily small value of $61e-$ 6. Note that the unimodal Gaussian is then equivalent to a standard deterministic policy trained with a mean squared error (MSE) loss. As commonly done [37], we stack a history of RGB frames into sequences of length four for our observations.
IV Experiments
[…]
Figure 3: Sequence of images from roll outs of place policies trained using data collected with LPP. The policies are able to place objects of varying properties in the scene using images from the wrist camera directly.
We use LPP to train robotic placement policies for two separate tasks: dishrask loading and table setting, as shown in Figure 3. In both cases, the place policy has to account for the weight and shape of each item, the initial grasp of the object and the state of the current scene to decide where and how to place (e.g., place angle). In this section, we present our quantitative results for the dishrack loading task as it better represents the type of task, with contact constraints from the environment, that LPP was designed for. We study the robustness of LPP in Section IV-A, the effect of noise augmentation on the performance of the learned placing policy in Section IV-B and the quality the policies trained with data from LPP and kinesthetic teaching in Section IV-C.
IV-A LPP Data Collection Robustness
@@
Figure 4: Tactile images from before and after a tactile regrasp (TR). Left: the plate’s edge is seen in the middle of the tactile image indicating a shallow grasp. Right: after a regrasp, the edge of the plate has shifted to a desired spot indicating a stable grasp.
Figure 5: Ablation study on compliant control for grasping (CCG) and tactile regrasp (TR). We measure the average number of episodes that can be collected autonomously based on the number of failures. Both CCG and TR play a crucial role in reducing the number of failures to zero, which allows us to collect the necessary amount of data seamlessly.
We require a robust data collection loop to collect a large amount of place demonstrations autonomously with LPP. We study the effects of compliant control for grasping (CCG) and tactile regrasp (TR), as introduced in Section III, on the robustness of LPP.
We observed two main causes that led to misplaced plates during data collection. First, the manipulator would stiffly grasp the object causing it to move and generate a large amount of force preload against the dishrack. During retrieval, significant object motion occurred at the moment when the object’s contact force with the environment was released. This unexpected shift in the object pose lead to less accurate placements since the reversal procedure no longer resulted in correct alignment between the object and the dishrack slot. With LPP, we mitigate this failure mode with CCG where we decrease the translational and rotational stiffness of the manipulator controller right before grasping the object. The second failure mode involved unstable grasps that lead to large relative motion between the object and the gripper. With LPP, we mitigate unstable grasps using TR to preemptively detect if a grasp will lead to relative object motion and to regrasp if needed. Based on empirical data, we find that unstable grasps typically occurred with shallow grasps. To this end, we design our regrasping module to recognize edges and calculate the relative change in EE pose required to recover a deep enough stable grasp.

### 20 Mar 2024 v2

@@ Abstract
We present placing via picking (PvP), a method to autonomously collect real-world demonstrations for a family of placing tasks in which objects must be manipulated to specific contact-constrained locations. With PvP, we approach the collection of robotic object placement demonstrations by reversing the grasping process and exploiting the inherent symmetry of the pick and place problems. Specifically, we obtain placing demonstrations from a set of grasp sequences of objects initially located at their target placement locations. Our system can collect hundreds of demonstrations in contact-constrained environments without human intervention by combining two modules: tactile regrasping and compliant control for grasps. We train a policy directly from visual observations through behavioral cloning, using the autonomously-collected demonstrations. By doing so, the policy can generalize to object placement scenarios outside of the training environment without privileged information (e.g., placing a plate picked up from a table). We validate our approach in home robotic scenarios that include dishwasher loading and table setting. Our approach yields robotic placing policies that outperform policies trained with kinesthetic teaching, both in terms of performance and data efficiency, while requiring no human supervision.
@@ I Introduction
Imitation learning (IL) provides an appealing, simple, and practical option by learning end-to-end control policies from expert demonstrations [1, 2]. The general approach of training expressive models on large and diverse datasets has proven to be quite effective outside of robotics. While there have been many attempts to transfer this paradigm to robotic systems [3, 4], robots have yet to achieve similar successes. A major bottleneck for IL-based approaches in these data regimes is the human time and effort required to collect the large amount of expert demonstrations in the physical world.
Having said this, many recent methods attempt to streamline [5, 6] or automate [7, 8, 9, 10] data collection. As highlighted by the authors of [9], the lack of robust and diverse autonomous data collection policies is a major limitation to scaling up IL. In this work, we are interested in investigating autonomous data collection for robotic object placement. As opposed to prior research, we tackle the full placing problem, without restricting our task to simple, small objects placed on flat, horizontal surfaces with no environment contact constraints. We assert that object and environment contact constraints can be a valuable, natural guide for learning how to place if handled properly.
We propose a novel approach named placing via picking (PvP) that automates the collection of expert demonstrations for a large subset of contact-constrained placing problems. We do so by taking advantage of a powerful grasp planner [11], tactile sensing, and compliant control. PvP is a self-supervised pipeline to autonomously collect expert placing demonstrations by leveraging the cyclical nature and inherent symmetry of the pick and place problems. Given an environment with objects initially positioned at their target place locations, we alternate between picking (i.e., grasping and retrieving) and placing, by time-reversing the retrieval trajectory. We generate demonstration data for object placement in a self-supervised manner—the picking phase provides the training supervision for the placing task. While appearing deceptively easy at first, we highlight the crucial importance of two modules, compliant control for grasping (CCG) and tactile regrasping (TR). CCG and TR enables (1) robust and uninterrupted pick-and-place and (2) leveraging environment contact-constraints as guides to inform object placement. LABEL:fig:title provides a visual summary of our approach. Our main contributions are:
@@
    PvP, a self-supervised data collection method for 6-DOF robotic object placements;
@@
    a compliant grasping and a tactile-driven regrasping strategy to achieve uninterrupted pick and place data collection without human intervention in contact-constrained environments;
@@
    a language-specified grasp planning pipeline; and
@@
    real-world experimental validation of PvP using collected demonstrations to train a vision-based policy capable of placing multiple plates in a dish rack and of setting a table.
@@ II Related Work
In this section, we review previous approaches to automate robot data collection in the real world and other work that generally leverage the idea of “working backwards”.
II-A Automatic Data Collection
A large number of existing policy learning approaches use prior policies to automate the data collection process. Examples include using trajectory optimization [12, 13] and, closer to our work, simple scripted pick-and-place policies [7, 8, 9] as forms of supervision. The authors of [7] use a scripted pick-and-place policy to bootstrap the initial dataset used for off-policy deep reinforcement learning with a real-world manipulator. In [9], the authors present AutoRT, a method that uses a combination of scripted, learned and tele-operated data collection policies to efficiently collect a large amount of language-specified manipulation demonstration data. Using a combination of a large language model (LLM) and visual language models (VLMs), AutoRT observes scenes in the wild and comes up with plausible tasks. Once a task is defined and chosen, AutoRT samples a data collection policy to attempt the task and collect real world data. In our work, we focus on developing a method for collecting demonstrations of placement with larger and more complex objects (i.e., objects that require a planner to grasp) in environments with potential contact constraints. Unlike prior work, we do not restrict the placing task to simpler small objects on flat horizontal surfaces. The authors of [9] mention the lack of robust and diverse real-world autonomous collection policies as a limitation of AutoRT; Our work is complementary with many of the prior frameworks and PvP could be used as a prior policy within them.
@@ II-B Working Backwards
The general idea of working backwards or time-reversal has been exploited in multiple contexts including generative modelling [14] and in learning visual predictive models for control [15, 16, 17]. Our self-supervised data collection method, which extends the concept of working backwards to the problem of robotic placement, is most similar in spirit to [18, 19, 20]. These methods consider objects that are assumed to be in their desired place configuration during data collection. The manipulator can then pick an object and move back to the initial goal location to generate data. However, we focus on placement as opposed to insertion as our task. Insertion policies generally only reason about local interactions, while placement involves reasoning over a larger scene with the potential for multiple, generally larger objects and place poses involved in the task. Furthermore, unlike prior work [18, 19, 20], our system does not require a human to guide the manipulator to the object and instead leverages tools from grasp planning. Our approach also resembles [21], where time-reversed trajectories are used to generate assembly data from disassembly data in a self-supervised manner. However, we focus on generating demonstration data for a closed-loop reactive policy that uses a 6-DOF pose-based action representation for placement. In contrast, the self-supervised data collection procedure introduced in [21] uses a 2D open loop pick-and-place action primitive.
II-C Robotic Manipulator Placing
Pick and place is a fundamental problem in robotic manipulation. A significant amount of research has focused on the picking problem [22, 23, 24], while placing is relatively less explored.
Previous approaches have explored the use of motion planning combined with place-specific objectives [25, 26]. For example, in [25], the authors demonstrate a hierarchical sampling-based motion planner that first finds poses which satisfy constraints (e.g., reachability) and then performs a local optimization for user-specified objectives (e.g., clearance). The place problem has also been studied from the perspective of geometric modelling [27], where the planar surfaces of a discretized object model are matched to planar patches in the environment. Importantly, 3D graphic models of the objects being placed as well as the environment are assumed to be available ahead of time.
Closer to our approach, other works have trained end-to-end place policies using reinforcement learning (RL) [28] and imitation learning (IL) [29]. In [29], the authors demonstrate an inverse reinforcement learning (IRL) technique capable of simultaneously learning a cost function and a control policy from human kinesthetic demonstrations of dish placing. The main contribution and novelty of our work is a data collection technique for robotic placing tasks that can be used in tandem with any methods that require expert demonstrations.
III Placing via Picking
We first present an overview of our self-supervised data collection method, PvP, in Section III-A and then describe a method for noise augmentation during data collection in Section III-D. Next, we provide details about the procedure used to train our IL policy in Section III-E.
@@ III-A Self-supervised Data Collection
Figure 1: The four steps involved in PvP, our autonomous demonstration data collection process for placing. We (a) generate grasps with an off-the-shelf grasp planner [11], (b) compliantly grasp the object to apply minimal forces to the environment while ensuring a stable grasp with tactile sensing, (c) retrieve the object with rotational compliance while storing the trajectory, and (d) generate placement demonstration data by rolling out the reversed grasp trajectories while storing the observations and actions.
PvP consists of a cycle with four main phases: (1) grasp planning, (2) grasping, (3) retrieving, and (4) placing by reversing; these steps are visualized in Figure 1. During training, the objects of interest are initially in their goal positions.
@@ III-A1 Grasp Planning
Figure 2: The steps involved in our language-driven grasp planning pipeline. We use (a) Grounding-Dino [30] for object bounding box detection based on text and (b) Segment Anything [31] on the cropped images to produce object specific masks. We then use (c) Contact-graspnet [11] for grasp generation on only the segmented objects (i.e., masked areas). Top: using “green plate” and “blue plate” as the targets for data collection. Bottom: using “bowl” and “mug” as the targets for data collection.
At the start of the data collection process, the manipulator moves to a pre-determined pose where the camera has a clear view of the objects of interest in the environment, as shown in Figure 1. We use the grasp planner Contact-GraspNet [11] to generate $L$ candidate grasp poses $\{\mathbf{T}_{g,i}\}_{i=0}^{L}$, where $\mathbf{T}_{g,i}\in SE(3)$ for all $i\in\{0,...,L\}$. Given a set of text-based descriptions of the objects of interest (e.g., “plates”, “green plate”, “cup” or, more generally, “objects”), potentially from a user or high-level LLM planner, we use Grounding-Dino [30] to find the object bounding boxes. Finally, we run Segment Anything [31] on each of the previously generated image crops from the bounding boxes to generate per-object masks. Contact-GraspNet can then use these masks to filter grasps that are not on objects of interest and keep only a pruned set of grasps $\{\mathbf{T}_{g,i}\}_{i=0}^{K}$, where $K\leq L$. We visualize an example of this process in Figure 2 for two different scenes. After pruning, we randomly select a single grasp pose, $\mathbf{T}_{\text{grasp}}$, from $\{\mathbf{T}_{g,i}\}_{i=0}^{K}$. To approach the object of interest without collisions, we include a pregrasp pose $\mathbf{T}_{\text{pregrasp}}$ that is defined with a translational offset above the grasp pose.
@@ III-A2 Grasping
Given the computed pregrasp and grasp poses, we move to each pose sequentially by linearly and spherically linearly interpolating to produce a smooth trajectory. An example of a grasp sequence is shown in Figure 1. During grasping, two modules play a critical role in the robustness of PvP: compliant control for grasping (CCG) and tactile regrasping (TR).
III-B Compliant Control for Grasping
@@
We use a cartesian impedance controller for manipulation. This choice allows us to set the level of compliance of the manipulator at various stages as needed. Before closing the gripper to grasp, we set the translational and rotational stiffness of the manipulator’s controller to minimal values. The low compliance of the robot arm then allows the manipulator and gripper to comply with the contact constraints and minimize its applied forces against the environment. Larger contact-constrained objects generate unnecessary reaction forces when gripped without compliance. These large forces trigger emergency stops and can damage the environment, and the robot. Specific to PvP, large reaction forces also disrupt and move the object relative to the gripper ultimately leading to inaccurate placements. On the other hand, when grasping with compliance, the environment contact constraints guide the manipulator and gripper towards a natural object placement pose.
@@
Figure 3: Tactile images from before and after a tactile regrasp (TR). Left: the plate’s contact surface area fills half of the tactile image indicating a shallow grasp. Right: after a regrasp, the contact surface area has increased indicating a stable grasp.